# NSSI-Net: Multi-Concept Generative Adversarial Network for Non-Suicidal Self-Injury Detection Using High-Dimensional EEG Signals in a Semi-Supervised Learning Framework

### v1

@@ VI-A Ablation Study
In this section, we conduct an ablation study to evaluate the contribution of each discriminator in the proposed model. Table II presents the model results using semi-supervised cross-subject ten-fold cross-validation under different discriminator configurations. (1) Signal-Specific Discriminator Contribution. Removing the signal-specific discriminator reduces the accuracy to 65.47±10.22, highlighting its crucial role in extracting relevant signal-based features. (2) Gender-Specific Discriminator Contribution. Without the gender-specific discriminator, the accuracy drops slightly to 66.63±09.66, indicating that gender-specific features contribute to overall classification performance, though less significantly than signal-specific features. (3) Domain-Specific Discriminator Contribution. When the domain-specific discriminator is removed, the accuracy decreases to 64.30±09.88, showing the importance of domain adaptation in improving the model’s performance. (4) Comparison with Traditional Domain Discriminator. Traditional domain adaptation methods are based on the DANN framework [48]. In line with previous approaches [53, 54], we treat both labeled and unlabeled source data as a unified domain for adaptation. Replacing the proposed discriminator setup with a traditional domain discriminator yields an accuracy of 67.17±12.99, which is lower than the proposed NSSI-Net’s performance of 70.00±13.90, underscoring the effectiveness of the proposed multi-concept discriminator framework. The results demonstrate the importance of each discriminator in enhancing the model’s performance, with the signal-specific and domain-specific discriminators having the most significant impact.
@@
TABLE II: The ablation study of the proposed NSSI-Net.
@@ VI-B Evaluating Model Performance under Varying Labeled Source Data Ratios in Semi-Supervised Learning
The availability of labeled source data during semi-supervised learning is crucial in determining the effectiveness of model training, especially when working with complex EEG signals. Understanding the impact of different ratios of labeled source data helps evaluate the trade-off between data quantity and model performance, providing insights into how much labeled data is required to achieve optimal outcomes. As shown in Fig. 5, we adjust the ratios of labeled source data from 5% to 85%.
The experimental results show that with only 5% of the labeled source domain data, the model struggles to learn effectively, achieving a relatively low accuracy rate of 50.97%. When the ratio increases to 10%, there is a significant improvement in accuracy, with the model reaching an average accuracy of 65.71%. This suggests that the model requires a minimum ratio of labeled source data to begin capturing meaningful patterns in the EEG signals. Beyond this 10% ratio, the accuracy continues to improve steadily, reaching a plateau when 75% is reached, with an accuracy of 70.00%. This suggests that the model, leveraging its semi-supervised learning framework, effectively utilizes both labeled and unlabeled data to maintain steady and robust performance, even when labeled source data is relatively limited.
@@ VI-C The Effect of Sampling Process
VI-D Topographic Analysis of Channel Contributions to NSSI Classification
@@
The normalized accuracy values are visualized in a topographic analysis, which allows us to discern spatial patterns of channel importance. As shown in Fig. 7, the parietal, frontal and occipital regions display notably higher significance in predicting NSSI behaviors compared to other brain areas. These regions are known to play key roles in emotional processing [55, 56, 57], visuospatial functions [58, 59], and self-referential thoughts [60, 61, 62], which could be critical in the development and maintenance of self-injurious behaviors.
@@
Figure 7: The topographic map shows the contribution of each EEG channel to the model’s performance. The color bar indicates the importance of each channel in distinguishing between DN+ and DN- groups. Red represents higher importance, while blue indicates lower contribution.
VI-E Conclusion
@@
In this paper, we introduce a novel semi-supervised framework with a multi-concept discriminator (NSSI-Net), designed for the detection of NSSI using EEG signals. NSSI-Net leverages spatial-temporal EEG patterns through a hybrid CNN-BiGRU encoder-decoder structure, which allows us to characterize the complexity of neural activities in adolescents with depression. The multi-concept discriminator is proposed to address the heterogeneous and complex nature of EEG data by focusing on multiple discriminative perspectives. It specifically focuses on capturing signal-specific, gender-specific, domain-specific, and disease-specific characteristics, enhancing the model’s ability to accurately classify NSSI-related features. Based on a self-collected dataset of 114 adolescents diagnosed with depression, the performance of the proposed NSSI-Net is validated using a semi-supervised cross-subject cross-validation approach. Comprehensive analysis of model components and hyperparameters are carefully conducted to assess the framework’s robustness. The results show that NSSI-Net achieves high efficiency and accuracy in detecting NSSI and provides interpretable insights into the neural mechanisms associated with NSSI. Future work would focus on enhancing the model’s ability to handle class imbalances and sample size issues, ensuring consistent performance across a broader range of real-world applications.
VII Acknowledgments
This work was supported in part by the National Natural Science Foundation of China under Grant 62276169, in part by the Medical-Engineering Interdisciplinary Research Foundation of Shenzhen University under Grant 2024YG008, in part by the Shenzhen University-Lingnan University Joint Research Programme, in part by the Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions under Grant 2022SHIBS0003, and in part by the STI 2030-Major Projects 2021ZD0200500.

### v2

@@ VI-A Ablation Study
In this section, we conduct an ablation study to evaluate the contribution of each discriminator in the proposed model. Table III presents the model results using semi-supervised cross-subject ten-fold cross-validation under different discriminator configurations. (1) Signal-Specific Discriminator Contribution. Removing the signal-specific discriminator reduces the accuracy to 65.47±10.22, highlighting its crucial role in extracting relevant signal-based features. (2) Gender-Specific Discriminator Contribution. Without the gender-specific discriminator, the accuracy drops slightly to 66.63±09.66, indicating that gender-specific features contribute to overall classification performance, though less significantly than signal-specific features. (3) Domain-Specific Discriminator Contribution. When the domain-specific discriminator is removed, the accuracy decreases to 64.30±09.88, showing the importance of domain adaptation in improving the model’s performance. (4) Comparison with Traditional Domain Discriminator. Traditional domain adaptation methods are based on the DANN framework [52]. In line with previous approaches [54, 55], we treat both labeled and unlabeled source data as a unified domain for adaptation. Replacing the proposed discriminator setup with a traditional domain discriminator yields an accuracy of 67.17±12.99, which is lower than the proposed NSSI-Net’s performance of 70.00±13.90, underscoring the effectiveness of the proposed multi-concept discriminator framework. The results demonstrate the importance of each discriminator in enhancing the model’s performance, with the signal-specific and domain-specific discriminators having the most significant impact.
@@
TABLE III: The ablation study of the proposed NSSI-Net.
@@ VI-B Evaluating Model Performance under Varying Labeled Source Data Ratios in Semi-Supervised Learning
The availability of labeled source data during semi-supervised learning is crucial in determining the effectiveness of model training, especially when working with complex EEG signals. Understanding the impact of different ratios of labeled source data helps evaluate the trade-off between data quantity and model performance, providing insights into how much labeled data is required to achieve optimal outcomes. As shown in Fig. 5, we adjust the ratios of labeled source data $\tau\%$ from 5% to 85%.
The experimental results show that with only 5% of the labeled source domain data, the model struggles to learn effectively, achieving a relatively low accuracy rate of 50.97%. When the ratio increases to 10%, there is a significant improvement in accuracy, with the model reaching an average accuracy of 65.71%. This suggests that the model requires a minimum ratio of labeled source data to begin capturing meaningful patterns in the EEG signals. Beyond this 10% ratio, the accuracy continues to improve steadily, reaching a plateau when 75% is reached, with an accuracy of 70.00%. This suggests that the model, leveraging its semi-supervised learning framework, effectively utilizes both labeled and unlabeled data to maintain steady and robust performance, even when labeled source data is relatively limited. However, when the labeled data ratio reaches 85%, a slight decrease in accuracy is observed. This is likely due to increased variability in the labeled data, introducing inconsistencies such as annotation noise or domain shifts. As a result, the model becomes more sensitive to these variations and leads to a decline in generalization ability. This emphasizes the importance of balancing labeled and unlabeled data in semi-supervised learning.
@@ VI-C The Effect of Sampling Process
VI-D The Effect of Parallel Processing in Feature Extraction
To compare the effectiveness of sequential and parallel feature extraction, we adjust the feature extraction pipeline. For clarity, we refer to the model where spatial features are extracted first, followed by temporal features, as the sequential model (the original architecture). Meanwhile, we refer to the model where spatial and temporal features are extracted simultaneously as the parallel model. Unlike the sequential model, which maintains a symmetric CNN-BiGRU encoder and BiGRU-CNN decoder, the parallel model lacks this symmetry, requiring an alternative decoding approach with a multi-layer convolutional decoder instead. To assess model performance, we conduct cross-validation under both processing strategies. The results indicate a significant performance decline from 70.00$\pm$13.90 (sequential model) to 58.03$\pm$5.59 (parallel model). This degradation may be primarily due to compromised temporal consistency. In the sequential model, BiGRU receives structured spatial-temporal features, which facilitates effective sequential modeling. In contrast, the parallel model forces BiGRU to process raw EEG data directly. This reduces its ability to capture long-term dependencies. These findings suggest that the original sequential structure remains superior for EEG-based NSSI detection, as it better preserves hierarchical feature integration and temporal dependencies.
VI-E Topographic Analysis of Channel Contributions to NSSI Classification
@@
The normalized accuracy values are visualized in a topographic analysis, which allows us to discern spatial patterns of channel importance. As shown in the Fig. 7 (a), the parietal, frontal and occipital regions demonstrate significantly higher importance in predicting NSSI behaviors compared to other brain areas. These regions are known to play key roles in emotional processing [56, 57, 58], visuospatial functions [59, 60], and self-referential thoughts [61, 62, 63], which could be critical in the development and maintenance of self-injurious behaviors.
When analyzing the gender-specific results, we observe the differences between the female and male subgroups. In female subgroup (Fig. 7 (b)), emotional processing engages both hemispheres, with bilateral activation in key regions, suggesting a more integrated neural processing approach. In contrast, the male subgroup rely more on unilateral processing (Fig. 7 (c)), with emotional responses exhibiting greater lateralized activity. These findings indicate gender-specific differences in emotional processing, with females utilizing both hemispheres and males showing more lateralized brain activity [64].
@@
Figure 7: The topographic maps show the contribution of each EEG channel to the model’s performance across three datasets: (a) Entire Dataset, (b) Female Subgroup, and (c) Male Subgroup. The color bar indicates the importance of each channel in distinguishing between the DN+ and DN- groups. Red represents higher importance, while blue indicates lower contribution.
VI-F Model Stability and Convergence Performance
To assess stability and convergence, we evaluate the overall performance of the designed multi-concept discriminator and the individual performance of each discriminator by examining their respective loss curves throughout the training process. We compute the average loss curves across all cross-validation folds to ensure a comprehensive evaluation of training dynamics, with the results presented in Fig. 8. For the overall performance of the multi-concept discriminator, the total loss curve (yellow) exhibits smooth and stable convergence after approximately 60 epochs. It indicates that the model successfully reaches an equilibrium where further training provides minimal gains and ensures strong generalization performance. For individual discriminators, the signal-specific discriminator’s loss curve (blue) stabilizes rapidly, converging around the 15th epoch. It suggests its effectiveness in distinguishing between true and false signals early in training. The gender- (green) and domain-specific (red) discriminators exhibit mild oscillations in the initial training phase, a common characteristic of adversarial networks. Nevertheless, both show a steady decline in loss after approximately 60 epochs, indicating effective feature learning. Similarly, the disease-specific discriminator (purple) stabilizes early, similar to the signal-specific discriminator, highlighting its strong feature extraction capability.
Furthermore, we conduct comprehensive experiments with various weight configurations for the hyperparameters ($\alpha$, $\beta$, $\delta$, $\theta$) used in Eq. 7. Specifically, we evaluate these hyperparameters ($\alpha$, $\beta$, $\delta$, $\theta$) with the following weight ratios: 1:1:1:1, 1:1:1:2, 1:1:2:1, 1:2:1:1, and 2:1:1:1. The corresponding results are 70.00$\pm$13.90, 65.12$\pm$09.07, 69.10$\pm$10.18, 69.11$\pm$13.93, and 63.56$\pm$08.73, respectively. These experiments reveal that the model achieves optimal performance with a balanced configuration ($\alpha$:$\beta$:$\delta$:$\theta$ = 1:1:1:1). It is noted that, increasing the weights of the disease-specific and signal-specific components leads to a significant decrease in performance. This suggests that the disease-specific and signal-specific discriminators converge more quickly during training. When these components are given excessive weight, they would dominate the learning process, causing the model to focus too narrowly on certain aspects. This analysis highlights the importance of balanced weight configurations for achieving optimal performance across all discriminators.
Figure 8: The total loss (yellow) represents the sum of the individual losses, and the curves show the loss values for the signal-specific (blue), gender-specific (green), domain-specific (red), and disease-specific (purple) discriminators over 80 epochs.
VI-G Conclusion
@@
In this paper, we introduce a novel semi-supervised framework with a multi-concept discriminator (NSSI-Net), designed for the detection of NSSI using EEG signals. NSSI-Net leverages spatial-temporal EEG patterns through a hybrid CNN-BiGRU encoder-decoder structure, which allows us to characterize the complexity of neural activities in adolescents with depression. The multi-concept discriminator is proposed to address the heterogeneous and complex nature of EEG data by focusing on multiple discriminative perspectives. It specifically focuses on capturing signal-specific, gender-specific, domain-specific, and disease-specific characteristics, enhancing the model’s ability to accurately classify NSSI-related features. Based on a self-collected dataset of 114 adolescents diagnosed with depression, the performance of the proposed NSSI-Net is validated using a semi-supervised cross-subject cross-validation approach. Comprehensive analysis of model components and hyperparameters are carefully conducted to assess the framework’s robustness. The results show that NSSI-Net achieves high efficiency and accuracy in detecting NSSI and provides interpretable insights into the neural mechanisms associated with NSSI. Future work would focus on enhancing the model’s ability to handle class imbalances and sample size limitations, including increasing the male sample size and expanding the dataset to ensure consistent performance across a broader range of real-world applications.